# AlignFormer: Modality Matching Can Achieve Better Zero-shot Instruction-Following Speech-LLM

### v1

@@ V-C More Zero-shot Speech Translation Evaluations
We further evaluate the instruction-following speech-LLM on a more general speech translation setting using CoVoST2 and FLEURS dataset based on audio-first training setup. Since we used EN ASR data only, the evaluation would be conducted on EN-X direction. The results are reported in Table IV. We also include the results of GT + Phi3.1 in the table to show Phi3’s machine translation capability with the ASR ground-truth (GT). The results can serve as the upper-bound for the AlignFormer because the translation capability is from LLM. As shown in the table, we can again see the high IFR for AlignFormer across different scenarios while the instruction following capability for the baseline is not quite stable. The BLEU scores for AlignFormer show that the model is usable in some scenarios but there are still rooms for improvements compared to the upper-bound. The translation capability in LLM has not been fully evacuated yet.
@@ VI Conclusions
In this paper, we studied zero-shot instruction-following speech-LLM and proposed to use instruction following rate (IFR) to evaluate the performance of instruction-following speech-LLM. To calculate IFR, we offered a systematic perspective by providing choices and the answer format in the prompt for different tasks. The IFR can be calculated by detecting the pattern in the LLM response. In addition, we analyzed the effect of audio position in training to the instruction following capability. We found that audio-first training achieved higher IFR than the instruction-first training due to a higher exposure to more diversified instructions. More importantly, we proposed a novel adapter, AlignFormer, to connect speech encoder and LLM for better speech and text modality matching. By using only ASR training data, AlignFormer can perform speech translation (ST) and speech question answering (SQA) tasks in a zero-shot manner. On the zero-shot evaluation of these tasks, AlignFormer achieved nearly 100% IFR in the audio-first setting and improved IFR from 15% to 49% in the instruction-frist setting. On a more general ST evaluation, we achieved average BLEU scores of 11.67 and 15.03 on CoVoST2 and FLEURS en-X translation, respectively. Our results show that it is possible to align the speech into LLM embedding space. We believe that this study is a big step towards the perfect speech and text modality matching.

### v2

@@ V-C More Zero-shot Speech Translation Evaluations
[…]
We further evaluate the instruction-following speech-LLM on a more general ST setting using CoVoST2 and FLEURS dataset based on audio-first training setup. Since we used EN ASR data only, the evaluation would be conducted on EN-X direction. The results are reported in Table IV. We also include the results of GT + Phi3.1 in the table to show Phi3’s machine translation capability with the ASR ground-truth (GT). The results can serve as the upper-bound for the AlignFormer because the translation capability is from LLM. As shown in the table, we can again see the high IFR for AlignFormer across different scenarios while the instruction following capability for the baseline is not quite stable. The BLEU scores for AlignFormer show that the model is usable in some scenarios but there are still rooms for improvements compared to the upper-bound. The translation capability in LLM has not been fully evacuated yet.
@@ V-D Ablation Study of AlignFormer
Finally, we calculate the similarity between the text and speech embeddings on E1-audio-first + AlignFormer using the librispeech test-other data. The average cosine similarity on the all utterances is 0.12. which is not high. We also calculate the average cosine similarity when AlignFormer is not used, the number of which is 0.08. The analysis shows the potential to incorporate additional distance loss between the speech and text embeddings for better modality matching, which is left for future work.
@@ VI Conclusions
In this paper, we studied zero-shot instruction-following speech-LLM and proposed to use instruction following rate (IFR) to evaluate the performance of instruction-following speech-LLM. To calculate IFR, we offered a systematic perspective by providing choices and the answer format in the prompt for different tasks. The IFR can be calculated by detecting the pattern in the LLM response. In addition, we analyzed the effect of audio position in training to the instruction following capability. We found that audio-first training achieved higher IFR than the instruction-first training due to a higher exposure to more diversified instructions. More importantly, we proposed a novel adapter, AlignFormer, to connect speech encoder and LLM for better speech and text modality matching. By using only ASR training data, AlignFormer can perform speech translation (ST) and speech question answering (SQA) tasks in a zero-shot manner. On the zero-shot evaluation of these tasks, AlignFormer achieved nearly 100% IFR in the audio-first setting and improved IFR from 15% to 49% in the instruction-frist setting. On a more general ST evaluation, we achieved average BLEU scores of 11.67 and 15.03 on CoVoST2 and FLEURS en-X translation, respectively. Our results showed that it is possible to align the speech into LLM embedding space and presented the advantage of AlignFormer over CTC ASR + LLM cascades solution. We believe that this study is a big step towards the perfect speech and text modality matching.